OpenStudy (anonymous):

Enter the value of the question mark as a fraction. Use a slash ( / ) as the fraction bar. 3^sqrt(5^5)=5^?

OpenStudy (anonymous):

You need to use log to figure this out: 3^sqrt(5^5) = 5^? 3^sqrt(3125) = 5^? 3^55.902 = 5^? 4.698 x 10^26 = 5^? \[\log_{5} 4.698 * 10^{26} = ?\] ? = 38 4/25